# High-Fidelity SLAM Using Gaussian Splatting with Rendering-Guided Densification and Regularized Optimization

### v1

@@ V Experiments
In this section, we demonstrate through quantitative and qualitative experiments that our method achieves better reconstruction results than the current state-of-the-art neural implicit SLAM methods and concurrent work using Gaussian splatting. Furthermore, the ablation studies of section V-D show that our Gaussian densification strategy and regularization terms can effectively improve reconstruction quality.
@@ V-A Experiment Setup
Evaluation Metrics: We use ATE (Absolute Trajectory Error [cm]) to measure the tracking performance, which aligns the estimated trajectory and the ground truth trajectory and then computes the root mean square error. To assess the quality of color image rendering, we employ peak signal-to-noise ratio (PSNR [dB]), structural similarity index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). PSNR evaluates pixel-wise RGB discrepancies, SSIM gauges structural resemblance, while LPIPS quantifies perceptual image patch similarity through neural network-based analysis. Since one can construct the mesh via TSDF-Fusion [23] with rendered depth along the estimated trajectory, we use metric Depth L1 [cm] to evaluate the geometric reconstruction as in [8]. All metrics are calculated by rendering full-resolution images along the estimated trajectory.
Datasets: We conduct our experiments on two datasets: Replica [24] (a synthetic dataset) and TUM-RGBD [25] (a real-world dataset); the synthetic dataset provides noiseless RGB and depth images while real-world datasets provide noisy inputs.
Baselines: We evaluate our results against both state-of-the-art SLAM methods using neural implicit representation —NICE-SLAM [4], ESLAM [5], Point-SLAM [8]—and 3D Gaussian Splatting—SplaTAM [11] and GS-SLAM [12].
@@ V-B Reconstruction Performance
TABLE I: The SLAM performance on the Replica [24] dataset. The best results are highlighted by first, second, and third. normal-$\uparrow$ means larger is better while normal-$\downarrow$ means smaller is better. Our method achieves the best results in most metrics.
@@
NICE-SLAM [4] Neural PSNR [dB] $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
@@
GS-SLAM [12] PSNR [dB] $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
@@
Compared to neural RGBD SLAM and more recent Gaussian Splatting SLAM, our method shows better rendering performance across both dataset, on real-world and synthetic data, as shown in table I and table II (it should be noted that since SplaTAM and GS-SLAM did not report rendering results on this dataset in their paper, we did not include them in the table).
Our method achieves the best results in color image rendering, but the second results in depth rendering in the Replica dataset. To explain why our method does not consistently outperform Point-SLAM on depth rendering, one must note that, in neural RGBD SLAM, appearance and geometry are often represented and optimized separately. In Point-SLAM, even though the same point cloud is used as primitives, each point is attached with a separate color feature descriptor and a geometric feature descriptor; also, two distinct neural networks are responsible for decoding color and depth. Hence, when optimizing the map parameters, color rendering and depth rendering do not affect each other. However, since our method uses Gaussian Splatting, the same Gaussian parameters $\{\mathbf{\mu},\mathbf{s},\mathbf{r},o\}$ are optimized for both color and depth rendering, which can impact the depth estimation.
@@
Ground Truth Point-SLAM [8] SplaTAM [11] Ours
Figure 3: The rendering results on the Replica dataset. The second and forth rows are zoomed-in details of the colored squares. Compared to Point-SLAM [8], our method generates sharper results; compared to SplaTAM [11], our method doen not have floaters.
@@ V-C Tracking Performance
On the Replica dataset, thanks to the improved reconstruction capabilities of our method shown in the previous section, we achieve the best tracking performance (see ATE RMSE in table I). However, on TUM-RGBD (table III), our method does not always perform the best. We hypothesise that, compared to methods using neural networks, our method is more sensitive to “noise”, such as motion blur and exposure changes present in real data.
TABLE II: The rendering performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third.
@@
NICE-SLAM [4] PSNR [dB] $\uparrow$ 13.83 12.00 11.39 12.89
@@
TABLE III: The tracking performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third. The evaluation metric for the trajectory is the ATE [cm].
@@
NICE-SLAM [4] 4.26 4.99 34.49 3.87
ESLAM [5] 2.47 3.69 29.73 2.42
Point-SLAM [8] 4.34 4.54 30.92 3.48
GS-SLAM [12] 3.30 - - 6.60
SplaTAM [11] 3.35 6.54 11.13 5.16
@@ V-D Ablation Study
In this section, we conduct ablation studies to demonstrate the effectiveness of the regularization terms and densification. For the regularization, we consistently render the first frame of Room0 in the Replica dataset at every step during the mapping. The fig. 4 shows rendered results of frame0 after the SLAM processing 350 frames. In the absence of regularization, the generated rendered image often exhibits artifacts such as floaters at the edge. Conversely, our approach ensures a cleaner output, with reduced or eliminated artifacts.
@@
Quantitatively, we show the overall result on the Room0 in Replica dataset in table IV.
@@
With regularization terms, we can achieve better reconstruction and tracking accuracy. Additionly, we also test with and without color and depth rendering densification, and the result (see the second row in table IV) reveals that the color and depth rendering densification can help improve rendering quality.
@@ V-E Runtime
Point-SLAM [8] 31.50 19.05 0.63 7.62
SplaTAM [11] 75.25 91.66 3.01 5.50
@@
Table V reports the runtime results on Room0 in Replica dataset running with a V100 GPU. For each iteration, Point-SLAM achieves the least time consumption because only a part of pixels are used in the tracking and mapping (in Point-SLAM, 1000 pixels are used in the mapping and 200 pixels are used in the tracking [^1] https://github.com/eriksandstroem/Point-SLAM). However, such a sparse sampling needs more iterations to converge, resulting in the longest time consumption for mapping; in addition, the sparse sampling decreases the rendering quality. SplaTAM and our method render full-resolution images, which require slightly more time at each iteration, but converge faster. It should be noted that our method is faster than SplaTAM due to implementation differences, since we render color and depth in one single rasterization process, as opposed to two rasterization processes in SplaTAM.
VI Conclusions
@@
Despite the promising results we achieved in the experiments, our method has some limitations. As shown in table III, our method does not always achieve good tracking results in real-world datasets due to motion blur and varying exposure. As shown in fig. 5, though the rendered image is of higher quality since the mapping accumulates all previous frames, the blurry ground truth image makes tracking difficult. In addition, when evaluating the rendering quality, the fact that the ground truth image is blurry negatively affects the evaluation metric.
@@
To improve tracking accuracy, it is popular to use bundle adjustment to optimize camera poses and the map simultaneously. However, we experimentally found that simply adopting bundle adjustment as in BARF [26] in our experiments will only make the result worse. Motivated by [27], we plan to investigate loop closure detection and optimize the trajectory in a pose graph optimization (PGO) way in the future.
In addition to tracking accuracy, for future work, we will also investigate the following directions. First, how to improve efficiency to make it a “real-time” system. Second, motivated by [28, 29], we will investigate how to introduce high-level information, such as semantics, into the mapping.

### v2

@@ V Experiments
In this section, we demonstrate through quantitative and qualitative experiments that our method achieves better reconstruction results than the current state-of-the-art neural implicit SLAM methods and concurrent work using Gaussian splatting. Furthermore, the ablation studies of Sec. V-D show that our Gaussian densification strategy and regularization terms can effectively improve reconstruction quality.
@@ V-A Experiment Setup
Evaluation Metrics: We use ATE (Absolute Trajectory Error [cm]) to measure the tracking performance, which aligns the estimated trajectory and the ground truth trajectory and then computes the root mean square error. To assess the quality of color image rendering, we employ peak signal-to-noise ratio (PSNR [dB]), structural similarity index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). PSNR evaluates pixel-wise RGB discrepancies, SSIM gauges structural resemblance, while LPIPS quantifies perceptual image patch similarity through neural network-based analysis. Since one can construct the mesh via TSDF-Fusion [curless1996volumetric] with rendered depth along the estimated trajectory, we use metric Depth L1 [cm] to evaluate the geometric reconstruction as in [sandstrom2023point]. All metrics are calculated by rendering full-resolution images along the estimated trajectory after reading all frames.
Datasets: We conduct our experiments on two datasets: Replica [straub2019replica] (a synthetic dataset) and TUM-RGBD [sturm2012benchmark] (a real-world dataset); the synthetic dataset provides noiseless RGB and depth images while real-world datasets provide noisy inputs.
Baselines: We evaluate our results against both state-of-the-art SLAM methods using neural implicit representation —NICE-SLAM [zhu2022nice], ESLAM [johari2023eslam], Point-SLAM [sandstrom2023point]—and 3D Gaussian Splatting—SplaTAM [keetha2023splatam] and GS-SLAM [yan2023gs].
Implementation Details: We test all the methods on one V100 GPU with 24 GB memory. The hyper-parameters used in our method are as follows: $\tau_{\text{color}}=0.6$, $\tau_{\text{depth}}=0.1$, $\lambda_{\text{color}}=0.5$, $\lambda_{\text{depth}}=1.0$, $\lambda_{\text{SSIM}}=0.2$.
@@ V-B Reconstruction Performance
TABLE I: The SLAM performance on the Replica [straub2019replica] dataset. The best results are highlighted by first, second, and third. $\uparrow$ means larger is better while $\downarrow$ means smaller is better. Our method achieves the best results in most metrics.
@@
NICE-SLAM [zhu2022nice] Neural PSNR [dB] $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
@@
GS-SLAM [yan2023gs] PSNR [dB] $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
@@
Compared to neural RGBD SLAM and more recent Gaussian Splatting SLAM, our method shows better rendering performance across both datasets, on real-world and synthetic data, as shown in Tab. I and Tab. II. (It should be noted that since SplaTAM and GS-SLAM did not report rendering results on this dataset in their paper, we did not include them in the table.)
Our method achieves the best results in color image rendering, but the second results in depth rendering in the Replica dataset. To explain why our method does not consistently outperform Point-SLAM on depth rendering, one must note that, in neural RGBD SLAM, appearance and geometry are often represented and optimized separately. In Point-SLAM, even though the same point cloud is used as primitives, each point has a separate color feature descriptor and a geometric feature descriptor; also, two distinct neural networks are responsible for decoding color and depth. Hence, when optimizing the map parameters, color rendering and depth rendering do not affect each other. However, since our method uses Gaussian Splatting, the same Gaussian parameters $\{\mathbf{\mu},\mathbf{s},\mathbf{r},o\}$ are optimized for both color and depth rendering, which can impact the depth estimation.
@@
Ground Truth Point-SLAM [sandstrom2023point] SplaTAM [keetha2023splatam] Ours
Figure 3: The rendering results on the Replica dataset. The second and forth rows are zoomed-in details of the colored squares. Compared to Point-SLAM [sandstrom2023point], our method generates sharper results; compared to SplaTAM [keetha2023splatam], our method doen not have floaters.
@@ V-C Tracking Performance
On the Replica dataset, thanks to the improved reconstruction capabilities of our method shown in the previous section, we achieve the best tracking performance (see ATE RMSE in Tab. I). However, on TUM-RGBD (Tab. III), our method does not always perform the best. We hypothesise that, compared to methods using neural networks, our method is more sensitive to “noise”, such as motion blur and exposure changes present in real data.
TABLE II: The rendering performance on TUM-RGBD [sturm2012benchmark] dataset. The best results are highlighted by first, second, and third.
@@
NICE-SLAM [zhu2022nice] PSNR [dB] $\uparrow$ 13.83 12.00 11.39 12.89
@@
TABLE III: The tracking performance on TUM-RGBD [sturm2012benchmark] dataset. The best results are highlighted by first, second, and third. The evaluation metric for the trajectory is the ATE [cm].
@@
NICE-SLAM [zhu2022nice] 4.26 4.99 34.49 3.87
ESLAM [johari2023eslam] 2.47 3.69 29.73 2.42
Point-SLAM [sandstrom2023point] 4.34 4.54 30.92 3.48
GS-SLAM [yan2023gs] 3.30 - - 6.60
SplaTAM [keetha2023splatam] 3.35 6.54 11.13 5.16
@@ V-D Ablation Study
In this section, we conduct ablation studies to demonstrate the effectiveness of the regularization terms and densification. For the regularization, we consistently render the first frame of Room0 in the Replica dataset at every step during the mapping. Figure 4 shows rendered results of frame0 after processing 350 frames. In the absence of regularization, the generated image often exhibits artifacts such as floaters at the edge. Conversely, our approach ensures a cleaner output, with reduced or eliminated artifacts.
@@
We show the overall quantitative results on the Room0 in Replica dataset in Tab. IV.
@@
With regularization terms, we can achieve better reconstruction and tracking accuracy. Additionally, we test with and without color and depth rendering densification, and the result (see the second row in Tab. IV) reveals that the color and depth rendering densification can help improve rendering quality.
@@ V-E Runtime
Point-SLAM [sandstrom2023point] 31.50 19.05 0.63 7.62
SplaTAM [keetha2023splatam] 75.25 91.66 3.01 5.50
@@
Table V reports the runtime results on Room0 in Replica dataset running with a V100 GPU. For each iteration, Point-SLAM achieves the least time consumption because only a subset of pixels are used in the tracking and mapping (in Point-SLAM, 1000 pixels are used in the mapping and 200 pixels are used in the tracking [^1] https://github.com/eriksandstroem/Point-SLAM). However, such a sparse sampling needs more iterations to converge, resulting in the longest time consumption for mapping; in addition, the sparse sampling decreases the rendering quality. SplaTAM and our method render full-resolution images, which requires slightly more time at each iteration, but converges faster. It should be noted that our method is faster than SplaTAM due to implementation differences, since we render color and depth in one single rasterization process, as opposed to two rasterization processes in SplaTAM.
VI Summary and Future Work
@@
Despite the promising results we achieved in the experiments, our method has some limitations. As shown in Tab. III, our method does not always achieve good tracking results in real-world datasets due to motion blur and varying exposure. As shown in Fig. 5, though the rendered image is of higher quality since the mapping accumulates all previous frames, the blurry ground truth image makes tracking difficult. In addition, when evaluating the rendering quality, the fact that the ground truth image is blurry negatively affects the evaluation metric.
@@
[…]